### v1

@@ Abstract
Scheduling virtual machines (VMs) to hosts in cloud data centers dictates efficiency and is an NP-hard problem with incomplete information. Prior work improved VM scheduling with predicted VM lifetimes. Our work further improves lifetime-aware scheduling using repredictions with lifetime distributions vs. one-shot prediction. The approach repredicts and adjusts VM and host lifetimes when incorrect predictions emerge. We also present novel approaches for defragmentation and regular system maintenance, which are essential to our data center reliability and optimizations, and are unexplored in prior work. We show that repredictions deliver a fundamental advance in effectiveness over one-shot prediction.
We call our novel combination of distribution-based lifetime predictions and scheduling algorithms Lifetime Aware VM Allocation (LAVA). LAVA improves resource stranding and the number of empty hosts, which are critical for large VM scheduling, cloud system updates, and reducing dynamic energy consumption. Our approach runs in production within Google’s hyperscale cloud data centers, where it improves efficiency by decreasing stranded compute and memory resources by $\sim$3and $\sim$2respectively, and increases availability for large VMs and cloud system updates by increasing empty hosts by 2.3-9.2 pp in production. We also show a reduction in VM migrations for host defragmentation and maintenance. In addition to our fleet-wide production deployment, we perform simulation studies to characterize the design space and show that our algorithm significantly outperforms the state of the art lifetime-based scheduling approach.
@@
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
Cloud data centers run an increasing fraction of the world’s compute. Efficient resource use in data centers is thus critical, from both an economic and environmental perspective. In these Infrastructure-as-a-Service (IaaS) environments, VM scheduling to physical hosts determines a large portion of end-to-end efficiency. Poor assignment of VMs may strand host resources – remaining host resources are too small or too imbalanced to accommodate additional VMs. VM allocation also impacts the number of empty hosts, which impacts idle power optimizations, resource utilization, the ability to schedule large VMs (obtainability), and system maintenance (e.g., for security patches).
@@
Much prior work addresses workload allocation in data centers Schwarzkopf et al. (2013); Verma et al. (2015); Cortez et al. (2017); Gog et al. (2016); Grandl et al. (2016); Zhang et al. (2014), including leveraging workload predictions Delimitrou & Kozyrakis (2014b; 2013); Park et al. (2018). The closest related work predicts job lifetimes (time between start and exit) to assist the scheduler in making better placement decisions Yadwadkar et al. (2014); Ferguson et al. (2012); Tumanov et al. (2016); Cortez et al. (2017); Buchbinder et al. (2021). Most recently, Barbalho et al. Barbalho et al. (2023) show bin packing improvements using Lifetime Alignment (LA) scheduling. LA predicts the lifetime of a VM at creation time, and then treats the predicted lifetime as fixed. As a result, mispredictions may tie up an entire host and, over time, model mistakes accumulate. For example, assume a model with a 1% chance of predicting a long-lived VM (running for months) as short-lived. If, over time, the scheduler places 70 presumably short-lived VMs on a host, there is a 50% chance that one of them is long-lived and prevents the host from freeing up until that VM exits. Correspondingly, Barbalho et al. show that their best algorithm achieves maximum efficiency with perfect prediction accuracy and degrades rapidly as the accuracy decreases (Barbalho et al. (2023), Figure 8). They thus introduce an alternative algorithm that achieves lower efficiency but is more tolerant to mispredictions.
@@
Our work achieves both tolerance to mispredictions and high efficiency, by repredicting VM and host lifetimes to correct mispredictions and improve accuracy over time. We take inspiration from work in C++ memory allocation that increases robustness of lifetime-based allocation by reassessing predictions on the fly Maas et al. (2020). In addition to bin packing, we address significant challenges not explored by prior work, including host defragmentation, maintenance, and stranding, which are essential components of Google’s cloud system design and approach to efficiency.
@@
Our key contributions are A) VM lifetime reprediction and adaptation to mispredictions, such as when a VM outlives its initial prediction; B) a novel ML model of VM lifetime probability distributions, which captures hard-to-predict VM behavior and prediction uncertainty; and C) three novel lifetime repredicting scheduling algorithms: 1) Lifetime Aware Scoring (LAS) incorporates predictions into an existing scoring function and corrects mispredictions. 2) Lifetime Aware VM Allocation (LAVA) more fundamentally redesigns the scheduler around lifetimes. 3) Lifetime-Aware ReScheduling (LARS) exploits lifetimes to reduce VM disruptions during host defragmentation and maintenance.
@@
Our approach is deployed fleet-wide across a planet-scale cloud infrastructure at Google. We evaluate our approach in production and through simulations with production traces, and compare to prior work. We validate that our simulations have high fidelity by comparing against production behavior.
@@
  • This paper shows that our key insight – probability distribution models of VM lifetimes – creates opportunities for new co-designed algorithms that repredict VM lifetimes and actively respond to mispredictions.
@@
  • It shows the new LAS, LAVA, and LARS algorithms increase empty host availability, reduce stranding, and reduce VM disruptions.
@@
  • Using production data and extensive simulation studies, we show that these algorithms are efficient and robust, outperforming the state of the art (SOTA).
@@
Similar to Maas et al. Maas et al. (2020), we find on-the-fly adjustment to mispredictions, versus one-off predictions, opens up a fundamentally different class of algorithms and optimizations, in this case for VM scheduling, and perhaps beyond. Our production deployment shows the practicality of this approach, and of deploying ML in the lower layers of the systems infrastructure stack more broadly.
@@ 2 Background
This section overviews VM lifetime properties, our production system, and metrics. Next, it describes the SOTA Lifetime Alignment algorithm (LA) Barbalho et al. (2023), to which we compare qualitatively and quantitatively.
@@ 2.1 Characterizing VM Lifetimes in Production
In our production environment, most VMs are short-lived, but most core hours are consumed by long-running VMs, similar to the generational hypothesis in garbage collection Lieberman & Hewitt (1983); Ungar (1984). Figure 1 shows our VM lifetime distributions. While 88% of all VMs live for less than 1 hour, 98% of resources are consumed by VMs that live for 1 hour or more (measured in CPU cores $\times$ time occupied). Without lifetime-specific handling of VMs, each VM would introduce the same amount of fragmentation, although only the placement of a small fraction of them has an impact on the overall resource efficiency. Prioritizing the bin packing quality of all VMs equally is thus highly inefficient, and motivates lifetime-based VM scheduling.
A key element of our approach is to view VM lifetimes from the perspective of distributions (Cumulative Density Functions or CDFs), as opposed to averages. We identify that VM features useful for prediction do not fully determine VM lifetimes. The best a traditional model could do would thus be to predict the expected (or average) lifetime for each VM. Figure 2 shows the average lifetime may not be particularly meaningful when there are both long and short-lived VMs with the same features. We therefore model and predict the lifetime of VMs as probability distributions. Note that this type of approach has a long history in survival analysis Wang et al. (2019) and in areas such as storage systems Zhou & Maas (2021). However, our novelty is predicting distributions and co-designing VM scheduling algorithms to leverage them in detecting and updating mispredictions.
@@ 2.2 VM Allocation
Our fleet consists of data centers in many geographic regions that run a mix of first and third party workloads. Cloud workloads have distinct host pools for distinct VM families. VM families represent different products, such as VMs optimized for high performance or cost efficiency. Broadly, VM families fall into two categories. 1) Slice-of-hardware families assign each VM a fixed partition of a host’s resources, e.g., CPU cores, DRAM, SSD, and GPUs. 2) Dynamically sized families share unused resources based on VM and host usage patterns and thus the amount of virtual CPUs and memory of VMs may exceed the physical resources on a host. This paper focuses on two VM families, C2 and E2, representative of these categories.
A scheduling framework called Borg Verma et al. (2015) assigns VMs to hosts within pools. Borg maintains a global view of all VMs and hosts in a given pool. When a VM creation request arrives, the scheduler computes the set of feasible hosts, i.e., hosts that have sufficient resources and match any hard constraints. For each feasible host, it then computes a score that determines the preference between them. This score incorporates a number of different business goals, such as spreading tasks across failure domains and bin packing efficiency. This scheduling setup is common Barbalho et al. (2023); Hindman et al. (2011); Verma et al. (2015).
Borg’s Waste Minimization bin packing algorithm explicitly optimizes for producing empty shapes of resources in multiple dimensions (e.g., CPU $\times$ DRAM $\times$ SSD) that match anticipated workload patterns. This approach was deployed several years ago based on experiments showing it achieves higher obtainability for large VMs and improves resource efficiency compared to Best Fit, the previous scheduler, which was similar to prior work Barbalho et al. (2023). However, Borg’s scoring function poses challenges for incorporating VM lifetimes, because it uses lexicographic ordering – one scoring dimension is evaluated at a time, with tie-breakers resolved by the next-lower scoring function. Bin packing is among its final dimensions, ensuring that improving bin packing does not degrade other business objectives.
@@ 2.3 Constraints & Objectives
Empty Hosts. We measure the number of empty hosts in a pool as a percentage of the total number of hosts. Empty hosts are an important resource as they are required to schedule large VMs that consume most or all of the host’s resources. They also enable host maintenance (e.g., rolling out a kernel or microcode security patch) without disrupting VMs with live migrations. Additionally, Google may put empty hosts in low power mode or move empty hosts between pools and configure them differently to serve various capacity needs. As such, 1 pp (percentage point) of increase in empty hosts directly corresponds to 1% more capacity available for these use cases. Other metrics of bin packing quality include packing density Barbalho et al. (2023), and are equivalent.
Resource Stranding. This metric measures free resource shapes and how many future VMs may be scheduled in them. Periodic inflation simulations measure stranding. We take a representative mix of VMs and simulate scheduling as many of them as possible until capacity is exhausted. The remaining resources on hosts represent stranded resources that cannot fit new VMs. For example, a host may contain free memory but no free CPUs. This memory is thus stranded. The metric is similar to admission control scoring functions Sajal et al. (2023) and captures unusable capacity.
VM Live Migrations. We aim to minimize live migrations. Live migration and fragmentation are closely connected. For some VM families, periodic live migrations will defragment resources. Live migration and resource efficiency may thus be traded against one another. This trade-off is controlled by policy and predates our work. It represents a fundamental difference from some other clouds. Maintenance events often trigger live migrations as well.
@@ 2.4 Baseline: Lifetime Alignment (LA)
We compare against LA (the Lifetime Alignment algorithm) from Barbalho et al. Barbalho et al. (2023). While other work on lifetime-based VM scheduling predates it, LA supersedes it and, to our knowledge, is the only published approach deployed in production data centers. At a high level, our setup (Figure 3) resembles Barbalho et al. We use ML to predict the lifetimes of VMs and incorporate predictions into scheduling decisions. Note that they use Best Fit scoring and do not consider maintenance and defragmentation cases, contributions of our paper.
For every VM allocation request, LA runs a gradient-boosting model to predict its lifetime class, which is over an exponentially increasing time interval ($2^{i}$ to $2^{i+1}$). Their model run on special inference servers. The allocator then preferentially assigns VMs to a host with the same lifetime class, using Best Fit. The lifetime class of a host is defined dynamically: it is the longest remaining time of any VM on that host, based on the VM’s initial lifetime prediction. If there is no host of the same lifetime class, LA picks another suitable host. If no such host exists, it assigns the VM to a new (previously empty) host.
@@
While LA Barbalho et al. (2023) tackled optimizing packing density (empty hosts), we extend the state-of-the-art by reducing stranding and our LARS algorithm decreases the number of migrations needed due to maintenance and defragmentation by determining the migration order. Since we create more empty hosts via lifetime-aware scheduling, we reduce the need for defragmentation.
@@ 2.5 Lifetime-aware C++ Memory Allocation
We observe that VM scheduling is similar to explicit memory allocation in C/C++ and other programming languages. The memory allocator needs to place objects, that never move, on pages such that pages are either maximally used or become fully empty Maas et al. (2020). VM scheduling is similar. VMs correspond to objects, hosts to pages, stranding to internal fragmentation, and empty hosts to external fragmentation. A key difference is that VM allocation is multi-dimensional. It manages multiple resources, including CPUs, memory, and SSD, while memory allocation is one-dimensional (object size).
@@
We take inspiration from LLAMA, designing our models and allocators to detect and correct mispredictions. We cannot directly adopt LLAMA because VM allocation is multi-dimensional and bin packing is not the only business objective. We thus incorporate the key insights of LLAMA into new algorithms, LAS, LAVA, and LARS.
3 ML Model

### v2

@@ Abstract
Scheduling virtual machines (VMs) on hosts in cloud data centers dictates efficiency and is an NP-hard problem with incomplete information. Prior work improved VM scheduling with predicted VM lifetimes. Our work further improves lifetime-aware scheduling using repredictions with lifetime distributions versus one-shot prediction. Our approach repredicts and adjusts VM and host lifetimes when incorrect predictions emerge. We also present novel approaches for defragmentation and regular system maintenance, which are essential to our data center reliability and optimizations, and are not explored in prior work. We show repredictions deliver a fundamental advance in effectiveness over one-shot prediction.
We call our novel combination of distribution-based lifetime predictions and scheduling algorithms Lifetime Aware VM Allocation (LAVA). LAVA reduces resource stranding and increases the number of empty hosts, which are critical for large VM scheduling, cloud system updates, and reducing dynamic energy consumption. Our approach runs in production within Google’s hyperscale cloud data centers, where it improves efficiency by decreasing stranded compute and memory resources by $\sim$3and $\sim$2respectively. It increases empty hosts by 2.3-9.2 pp in production, reducing dynamic energy consumption, and increasing availability for large VMs and cloud system updates. We also show a reduction in VM migrations for host defragmentation and maintenance. In addition to our fleet-wide production deployment, we perform simulation studies to characterize the design space and show that our algorithm significantly outperforms the prior state of the art lifetime-based scheduling approach.
@@ 1 Introduction
Cloud data centers run an increasing fraction of the world’s compute. Efficient resource use in data centers is thus critical, from both an economic and environmental perspective. In Infrastructure-as-a-Service (IaaS) environments, VM scheduling to physical hosts determines a large portion of end-to-end efficiency. Poor assignment of VMs may strand host resources – making the remaining host resources too small or too imbalanced to accommodate additional VMs. VM allocation also impacts the number of empty hosts, which impacts idle power optimizations, resource utilization, the ability to provision large VMs (obtainability), and system maintenance velocity (e.g., for security patches).
@@
Much prior work addresses workload allocation in data centers Schwarzkopf et al. (2013); Verma et al. (2015); Cortez et al. (2017); Gog et al. (2016); Grandl et al. (2016); Zhang et al. (2014), including leveraging workload predictions Delimitrou & Kozyrakis (2014b; 2013); Park et al. (2018). The closest related work predicts job lifetimes (time between start and exit) to assist the scheduler in making better placement decisions Yadwadkar et al. (2014); Ferguson et al. (2012); Tumanov et al. (2016); Cortez et al. (2017); Buchbinder et al. (2021). Most recently, Barbalho et al. Barbalho et al. (2023) show bin packing improvements using Lifetime Alignment (LA) scheduling. LA predicts the lifetime of a VM at creation time, and then treats the predicted lifetime as fixed. As a result, mispredictions may tie up an entire host and, over time, model mistakes will accumulate. For example, assume a model with a 1% chance of predicting a long-lived VM (running for months) as short-lived. If, over time, the scheduler places 70 presumably short-lived VMs on a host, there is a greater than 50% chance that one of them is long-lived, which would prevent the host from freeing up until that VM exits. Correspondingly, Barbalho et al. show that their best algorithm achieves maximum efficiency with perfect prediction accuracy and degrades rapidly as the accuracy decreases (Barbalho et al. (2023), Figure 8). They thus introduce an alternative algorithm that achieves lower efficiency but is more tolerant to mispredictions.
@@
Our work achieves both tolerance to mispredictions and high efficiency by repredicting VM and host lifetimes to correct mispredictions and improve accuracy over time. We take inspiration from C++ memory allocation work that increases the robustness of lifetime-based allocation by reassessing predictions on the fly Maas et al. (2020). For example, if a VM outlives its predicted lifetime, the algorithms change which other VMs are preferred to be co-scheduled with it on the VM’s host. In addition to host bin packing, we address significant challenges not explored by prior work, including host defragmentation, maintenance, and stranding, which are essential components of Google’s cloud system design and approach to efficiency.
@@
Our key contributions are A) VM lifetime reprediction and adaptation to mispredictions, such as when a VM outlives its initial prediction; B) a novel ML model of VM lifetime probability distributions, which captures hard-to-predict VM behavior and prediction uncertainty; and C) three novel lifetime repredicting scheduling algorithms: 1) Non-Invasive Lifetime Aware Scheduling (NILAS) incorporates predictions into an existing scoring function. 2) Lifetime Aware VM Allocation (LAVA) more fundamentally redesigns the scheduler around lifetimes. 3) Lifetime-Aware ReScheduling (LARS) exploits lifetimes to reduce VM disruptions during host defragmentation and maintenance.
@@
We deployed our approach fleet-wide across Google’s planet-scale cloud infrastructure. We evaluate it in production and in high-fidelity simulations with production traces, and compare these results to prior work. Some of the key findings of our paper include the following:
@@
  • This paper shows how to exploit our key insight – using probability distribution models of VM lifetimes – to co-design algorithms that repredict VM lifetimes and actively respond to mispredictions.
  • While NILAS and the prior state-of-the-art (SOTA) LA algorithm place VMs with similar lifetime together, the key idea of LAVA is the opposite – it puts shorter-lived VMs on hosts with one or more longer-lived VM, creating hosts with a wider range of lifetimes. This approach avoids extending the time at which the longest-lived VM will exit and the host will be empty.
@@
  • We show the NILAS, LAVA, and LARS algorithms increase empty host availability (2.3-9.2 percentage points (pp)), reduce stranding ($\sim$2-3%), and reduce VM disruptions (4.5%). Note that a consistent 1 pp improvement in stranding or empty hosts can be equivalent to saving 1% of a cluster’s capacity.
@@
  • Using production data and extensive simulation studies, we show that these algorithms are efficient and robust, outperforming the prior SOTA.
@@
Similar to Maas et al. Maas et al. (2020), we find that on-the-fly adjustment to mispredictions, versus one-off predictions, opens up a fundamentally different class of algorithms and optimizations, in this case for VM scheduling, and perhaps beyond. Our production deployment shows the practicality of this approach, and of deploying ML in the lower layers of the systems infrastructure stack more broadly.
@@ 2 Background
We first give an overview of VM lifetime properties, our production system, and metrics. Next, we describe the prior SOTA Lifetime Alignment algorithm (LA) Barbalho et al. (2023), to which we compare qualitatively and quantitatively.
@@ 2.1 Characterizing VM Lifetimes in Production
In our production environment, most VMs are short-lived, but most core hours are consumed by long-running VMs, similar to the generational hypothesis in garbage collection Lieberman & Hewitt (1983); Ungar (1984). Figure 1 shows our VM lifetime distributions. While 88% of all VMs live for less than 1 hour, 98% of resources are consumed by VMs that live for 1 hour or more (measured in CPU cores $\times$ time occupied). Without lifetime-specific handling of VMs, each VM would introduce the same amount of fragmentation, although only the placement of a small fraction of them has an impact on the overall resource efficiency. Prioritizing the bin packing quality of all VMs equally is thus highly inefficient, which motivates lifetime-based VM scheduling.
A key element of our approach is to view VM lifetimes from the perspective of distributions (Cumulative Density Functions or CDFs), as opposed to just averages. We identify that VM features useful for prediction do not fully determine VM lifetimes. The best a traditional model could do would thus be to predict the expected (or average) lifetime for each VM. Figure 2 shows the average lifetime may not be particularly meaningful when there are both long and short-lived VMs with the same features. We therefore model and predict the lifetime of VMs as probability distributions. Note that this type of approach has a long history in survival analysis Wang et al. (2019) and in areas such as storage systems Zhou & Maas (2021). However, our novelty is predicting distributions and co-designing VM scheduling algorithms to leverage them in detecting and updating mispredictions.
@@ 2.2 VM Allocation
Our fleet consists of data centers in many geographic regions that run a mix of first and third party workloads. Cloud workloads have distinct host pools for distinct VM families. VM families represent different products, such as VMs optimized for high performance or cost efficiency. Broadly, VM families fall into two categories. 1) Performance-optimized, slice-of-hardware families assign each VM a fixed partition of a host’s resources, e.g., CPU cores, DRAM, SSD, and GPUs. 2) Dynamically sized families share unused resources based on VM and host usage patterns and thus the amount of virtual CPUs and memory of VMs may on occasion exceed the physical resources on a host. This paper focuses on the performance-optimized C2 family and cost-optimized E2 family, representative of these categories.
A scheduling framework called Borg Verma et al. (2015) assigns VMs to hosts. Borg maintains a global view of all VMs and hosts in a given host pool. When a VM creation request arrives, the scheduler computes the set of feasible hosts, i.e., hosts with sufficient resources that match any hard constraints. For each feasible host, it computes a score that determines the preference between them. This score incorporates a number of different business goals, such as spreading tasks across failure domains and bin packing efficiency. This scheduling setup is common Barbalho et al. (2023); Hindman et al. (2011); Verma et al. (2015).
Borg’s Waste Minimization bin packing algorithm explicitly optimizes for producing empty shapes of resources in multiple dimensions (e.g., CPU $\times$ DRAM $\times$ SSD) that match anticipated workload patterns. Borg deployed this approach several years ago because it achieves higher obtainability for large VMs and improves resource efficiency compared to Best Fit, the previous scheduler, which was similar to prior work Barbalho et al. (2023). However, Borg’s scoring function poses challenges for incorporating VM lifetimes, because it uses lexicographic ordering – one scoring dimension is evaluated at a time, with tie-breakers resolved by the next-lower scoring function. Bin packing is among its final dimensions, ensuring that improving bin packing does not degrade other business objectives.
@@ 2.3 Constraints & Objectives
Empty Hosts. We measure the percentage of hosts in a pool that is empty. Empty hosts are required to schedule large VMs that consume most or all of the host’s resources. They also increase host maintenance velocity when rolling out a kernel or microcode security patch: By increasing empty hosts, applying the update to empty hosts first, and preferring new VMs land on updated hosts, we speed up maintenance and reduce VM disruptions due to live migrations. Additionally, Google puts empty hosts in low power mode and moves empty hosts between pools, configuring them differently to serve various capacity needs. As such, 1 pp of increase in empty hosts directly corresponds to 1% more capacity available for these use cases. Other metrics of bin packing quality include packing density Barbalho et al. (2023), and are equivalent (Appendix D).
Resource Stranding. This metric measures free resource shapes and how many future VMs may be scheduled in them. Periodic inflation simulations measure stranding. We take a representative mix of VMs and simulate scheduling as many as possible until capacity is exhausted. The remaining resources on hosts represent stranded resources that cannot fit new VMs. For example, a host may contain free memory but no free CPUs. This memory is thus stranded. The metric is similar to admission control scoring functions Sajal et al. (2023) and captures unusable capacity when the zone is full.
VM Live Migrations. We aim to minimize live migrations. Live migration and fragmentation are closely connected. For some VM families, periodic live migrations will defragment resources. Live migration and resource efficiency may thus be traded against one another. This trade-off is controlled by policy and predates our work. It represents a fundamental difference from some other clouds. Maintenance events trigger live migrations as well.
@@ 2.4 Baseline: Lifetime Alignment (LA)
We compare against LA (the Lifetime Alignment algorithm) from Barbalho et al. Barbalho et al. (2023). While other work on lifetime-based VM scheduling predates it, LA supersedes this work and, to our knowledge, is the only published approach deployed in production data centers. At a high level, our setup (Figure 3) resembles Barbalho et al. We use machine learning to predict the lifetimes of VMs and incorporate predictions into scheduling decisions. Note that their work uses Best Fit scoring. They do not correct mispredictions nor do they consider maintenance and defragmentation, contributions of our paper.
For every VM allocation request, LA runs a gradient-boosting model to predict its lifetime class, which is over an exponentially increasing time interval ($2^{i}$ to $2^{i+1}$). Their models run on special inference servers. The allocator then preferentially assigns VMs to a host with the same lifetime class, using Best Fit. The lifetime class of a host is defined dynamically: it is the longest remaining time of any VM on that host, based on the VM’s initial lifetime prediction. If there is no host of the same lifetime class, LA picks another suitable host. If no such host exists, it assigns the VM to a new (previously empty) host.
@@
While LA Barbalho et al. (2023) tackled optimizing packing density (empty hosts), we extend the state-of-the-art by reducing stranding and our LARS algorithm decreases the number of migrations needed due to maintenance and defragmentation by determining the migration order. In addition, since we create more empty hosts via lifetime-aware scheduling, we further reduce the need for defragmentation.
@@ 2.5 Lifetime-aware C++ Memory Allocation
We observe that VM scheduling is similar to explicit memory allocation in C/C++ and other languages. The memory allocator needs to place objects, that never move, on pages such that pages are either maximally used or become empty Maas et al. (2020). VM scheduling is similar. VMs correspond to objects, hosts to pages, stranding to internal fragmentation, and empty hosts to external fragmentation. A key difference is that VM allocation is multi-dimensional. It manages multiple resources, including CPUs, memory, and SSD, while memory is one-dimensional (object size).
@@
We take inspiration from LLAMA, designing our models and allocators to detect and correct mispredictions. We cannot directly adopt LLAMA because VM allocation is multi-dimensional and bin packing is not the only business objective. We do, however, incorporate the key insights of LLAMA into new algorithms, NILAS, LAVA, and LARS.
3 ML Lifetime Prediction Model